OpenStudy (anonymous):

Find the roots of the polynomial equation. x^2 - 8x - 48 = 0 A. –6, 8 B. –4, 12 C. –12, 4 D. –8, 6

OpenStudy (anonymous):

(x-12)(x+4)=0

OpenStudy (anonymous):

C?

sam (.sam.):

From (x-12)(x+4)=0 You should set each bracket equals to zero and solve for x

OpenStudy (anonymous):

so wich one is it?

sam (.sam.):

x-12=0 and x+4=0

OpenStudy (anonymous):

c?